OpenStudy (anonymous):

D=4,a1=3 find nth term

OpenStudy (anonymous):

nth term = a + (n-1)d = 3 + (n-1)(4) = 3 + 4n - 4 = -1 + 4n nth term is -1 + 4n or 4n - 1

OpenStudy (anonymous):

Obviously this is an arithmetic sequence whose common difference, d , is 4.

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

Thanks for all of your help